what is the equation for this problem?
Find three consecutive integers such that the sum of the first and twice the third is equal to four times the second decreased by five?

Answer by macston(5194)

.
The phrasing is ambiguous
.
F=first integer; S=second integer=F+1; T=third integer=F+2
.
F+2T=4(S-5) or F+2T=4S-5
.
F+2(F+2)=4(F+1-5) or F+2(F+2)=4(F+1)-5
.
F+2F+4=4F-16 or F+2F+4=4F+4-5
3F+4=4F-16 or 3F+4=4F-1
20=F or 5=F
S=21 or 6
T=22 or 7
.
ANSWER: The numbers are 5,6, and 7 or 20,21, and 22.